# Confronting Dark Matter Capture Rate with Continuous Gravitational Wave Probe of Local Neutron Stars

### v2

@@ IV Current and Future generation telescopes
Combining data from gamma-ray and neutrino observations enables us to construct a comprehensive picture of DM distribution and behavior across different cosmic scales. Moreover, cross-correlating these observations with other astronomical datasets, such as those from gravitational wave detectors and galaxy surveys, enriches our understanding of DM’s role in shaping the cosmos. Multiwavelength studies thus stand as a powerful strategy to decipher the enigmatic properties of DM and elucidate its profound impact on the Universe.
In sec. III.3, we discuss the DM annihilation spectra through long-lived mediators. In this section, we probe the expected gamma-ray and neutrino flux resulting from DM captured inside NS with the current and future generations of neutrino and gamma-ray telescopes.
@@ IV.1 With current and future neutrino telescopes
Neutrino telescopes offer a novel avenue for probing DM Abbasi et al. (2022a); Miranda et al. (2023) by detecting neutrinos. These telescopes complement traditional detection methods, providing insights into DM distribution and properties through neutrino emissions from potential DM sources across the Universe.
IceCube, a massive cubic-kilometer neutrino detector situated at the South Pole, was completed in 2010. Originally proposed based on theoretical calculations in 1998, it was envisioned to be sensitive enough to detect high-energy neutrinos from phenomena such as AGN jets or GRBs Waxman and Bahcall (1999). As the first experiment of its kind, IceCube has achieved significant milestones, including the discovery of a diffuse extraterrestrial neutrino flux in 2013 Aartsen et al. (2013) and the detection of neutrino emission from a flaring blazar in 2017 Aartsen et al. (2018a, b). Despite dedicated efforts to identify the sources of cosmic neutrinos through various analyses, such as all-sky survey Aartsen et al. (2020a), transient Abbasi et al. (2022b); Aartsen et al. (2018c, 2020b), and AGN catalog stacking searches Abbasi et al. (2022c); Aartsen et al. (2017a), conclusive results remain elusive. This suggests the possibility of multiple weaker sources contributing to the diffuse flux Murase et al. (2016), such as starburst galaxies Ambrosone et al. (2021), which would necessitate improved pointing resolution for resolution Fang et al. (2016).
IceCube-Gen2, an extension of the IceCube Neutrino Observatory in Antarctica, aims to explore the high-energy neutrino sky from TeV to PeV energies Aartsen et al. (2019); Abbasi et al. (2023). It promises five times better sensitivity for neutrino sources compared to the current IceCube detector. With 120 new strings and larger spacing between them, IceCube-Gen2 will enhance sensitivity for neutrinos above 10 TeV and target detection of neutrinos exceeding 100 PeV Abbasi et al. (2023). The ongoing IceCube upgrade will improve detection thresholds down to 1 GeV, enhancing capabilities for various physics studies Abbasi et al. (2021). IceCube-Gen2’s larger instrumented area will enable the detection of fainter neutrino sources, positioning it as the world’s leading neutrino observatory. It holds promise for diverse scientific inquiries, including DM searches, particle physics studies, and the detection of supernova neutrinos Ishihara (2021).
@@ IV.2 With current and future gamma-ray telescopes
Multi-messenger astrophysics has become a groundbreaking reality, propelled by recent advancements in gravitational wave, gamma-ray astronomy and high-energy neutrino detection. These breakthroughs have ushered in a new era of understanding the high-energy astrophysical universe and the mechanisms underlying its most energetic phenomena.
@@
As the next-generation gamma-ray telescope to observe very- to ultra-high-energy gamma rays, the Southern Wide-field Gamma-ray Observatory (SWGO), a water Cherenkov telescope array, stands at the forefront of this revolutionConceição (2023). Spanning energies from about 30 GeV to a few PeV, SWGO provides a wide field and high-duty cycle view of the southern sky Conceição (2023), complementing existing particle arrays in the Northern Hemisphere like Large High Altitude Air Shower Observatory (LHAASO) LHA (2021). Speaking of which, LHAASO, is a multi-purpose facility designed to study cosmic rays and gamma rays across a broad energy spectrum, ranging from sub-TeV to beyond 1 PeV Cao (2010).
@@ V.1 Differential flux sensitivity
Finally, for TRIDENT which is expected to be the most sensitive neutrino telescope in the upcoming days, we use the 90% C.L. median sensitivity for the point-like sources from the expected all-sky survey Ye et al. (2022). In this context, we want to mention that for all neutrino telescopes, we only consider their sensitivity from track-like events for the sources near declination, $\delta=0^{0}$.
For the gamma-ray telescopes, such as CTA (cta ; Donini et al. (2023)), SWGO (Conceição (2023)), LHAASO (LHA (2021); Conceição (2023)) and Fermi-LAT (fer (2021); Ajello et al. (2021)), we use the differential flux limits for point-like sources from their respective performance guidelines.
Figure 1: Differential flux ($E_{\nu,\gamma}^{2}~{}d\phi_{\nu,\gamma}/dE_{\nu,\gamma}$) sensitivity of all our selected detectors.
@@ V.2 Variation of differential flux sensitivity of neutrino telescopes with declination
In this section, we study the neutrino signals from NSs resulting from the decay of long-lived mediators produced from DM annihilation in the light of current and future-generation neutrino telescopes. This study specifically focuses on the DM-nucleon scattering cross-section in equilibrium, aiming to explore whether NSs offer a promising avenue for probing small DM-nucleon cross-sections. Throughout our investigation, we assume that DM number abundance has reached equilibrium, given the age of neutron stars considered to be larger ($\geq 1$ Myr) than $t_{\rm eq}$. All of our selected neutrino telescopes are equally sensitive for both spin-independent (SI) and spin-dependent (SD) DM-nucleon scattering cross-sections as NSs primarily consist of neutrons.
Figure 2: Variation of flux with declination for Box-like spectra for two DM mass values.
@@
Fig. 2 indicates that for IceCube Aartsen et al. (2020a) and IceCube Gen-2Omeliukh et al. (2021), the highest sensitivity is concentrated around the equator, corresponding to optimal discovery potential, while it is weakest for declination values near sin($\delta$) = ±$\pm 1$ Aartsen et al. (2020a). KM3NeT (with 6 years of projected data) Aiello et al. (2019) and TRIDENT Ye et al. (2022), provide equal sensitivity for all the regions of the sky. This underscores the challenge of IceCube and IceCube-Gen2 of detecting events of interest amidst the background, particularly in regions where a stronger signal is required. For a source located in the southern sky, TRIDENT will have nearly 4 orders of magnitude improvement in sensitivity Ye et al. (2022) compared to IceCube.
Consequently, given the present sensitivity of all these neutrino detectors, TRIDENT has the potential to detect the local NSs for DM mass around 1 GeV with $\sigma_{\chi n}$ = $6~{}\times~{}10^{-46}$ cm-2. Whereas for DM mass for $m_{\chi}~{}\sim$ 5 GeV, as shown in Fig. 2, the current and future telescopes are not sufficient unless brighter sources are present or DM-nucleon scattering is enhanced further. The choice of $\sigma_{\chi n}$ = $6~{}\times~{}10^{-46}$ cm-2 is motivated by its proximity to the saturation cross-section of our designated local NSs. Once surpassing this threshold, multi-scattering becomes a factor to be taken into account.
@@ V.3 Expected bounds on $\sigma_{\chi n}$ from local NSs
In Fig. 3, with conservative measures, we translate the differential flux sensitivity (Fig. 1) of our selected detectors and derive the stacked constraints on $\sigma_{\chi n}$ as a function of the DM mass, $m_{\chi}$ by combining the flux limits from 8 local NSs (Eqs. 7 and 8). The limits shown in Fig. 3 emerge from the intricate interplay of the DM capture rate, DM density profile, and DM mass range and serve as the current constraints on $\sigma_{\chi n}$ for 8 nearby local NSs. We derive the stacked limits for box spectra (long-lived mediators) for the NFW density profile. To ensure the robustness of our findings, we validate the results against those obtained for the cored Burkert profile Burkert (1995); Salucci and Burkert (2000). However, our analysis reveals $\sigma_{\chi n}$ has no significant dependence on the density profile for local NSs. Consequently, this paper exclusively features results derived from the NFW halo profile.
Fig.3 highlights that with 8 local NSs, we only obtain the limits for DM mass $<$ 1 GeV, and IceCube-Gen2, IceCube, TRIDENT can provide any meaningful bounds. We do not obtain any bounds for gamma-ray detectors. Additionally, in Fig. 3, we further compare the limits with experimental bounds obtained from direct detection bounds, such as CDEX Jiang et al. (2018), PICO-60 Amole et al. (2019), and XENON1T Aprile et al. (2019) and astrophysics limits obtained from Sun performed by IceCube DeepCore Abbasi et al. (2022a). For DM mass $<$ 1 GeV, a handful of nearby NSs provide stronger bounds than obtained from XENON1T Aprile et al. (2019). Therefore, Fig. 3 signifies that with the current limits from 8 local NSs, we only derive the bounds for DM mass $<$ 1 GeV.
@@
We now derive the bounds on $\sigma_{\chi n}$ from the cumulative emission of the local NS population within 0.6 kpc from Earth. Following the population model proposed by ref. Reed et al. (2021), we anticipate $N_{\rm ns}~{}\approx~{}10^{5}$ within 0.6 kpc from earth for NS ellipticity $\epsilon~{}\sim~{}10^{-5}$. Using the local NS population in the vicinity of Earth from CGW search, we estimate the projected bounds for gamma-ray and neutrino telescopes from the cumulative emission coming from around 600000 NSs.
In Fig. 4, we show the DM annihilation spectra (for both gamma-ray and neutrino) through long-lived mediators for energy ranging between $0.01-10^{8}$ GeV for three scattering cross-section ($\sigma_{\chi n}$) values. As mentioned before, here we consider the expected number of local NSs $N_{\rm ns}\approx 10^{5}$, probed by CGW within 0.6 kpc surrounding Earth. The DM-nucleon scattering cross-section, $\sigma_{\chi n}~{}=10^{-46}~{}{\rm{cm}}^{2}$, which is very close to the saturation cross-section limits, produces the weakest flux. The fluxes get stronger with increasing the $\sigma_{\chi n}$ values, which also motivates the multiple scatterings, and they can be probed with our selected detectors both for gamma-ray and neutrino.
@@
Figure 4: DM annihilation spectra through long-lived mediators for three scattering cross-section values.
In Fig. 5, we show the projected bounds for the local NS population for box-like spectra. This figure additionally signifies that we can expect the strongest bounds from TRIDENT (for neutrino) for DM mass $<10^{3}$ GeV and SWGO (for gamma-ray) for DM mass $>10^{3}$ GeV. Respective bounds on $\sigma_{\chi n}$ for larger DM mass ($10^{3}-10^{5}$ GeV), resulting from the analysis on CTA, IceCube-Gen2, LHAASO observation is stronger than the limits from IceCube results with Sun and comparable with indirect search from H.E.S.S. based on galactic NS population.
@@
Figure 5: Comprehensive limits on $\sigma_{\chi n}$ from local NS population with CGW searches (solid lines) and comparison with other DM direct and indirect bounds (dashed lines). See text for details.
In Fig. 5, we extend our analysis to compare our findings with the results obtained from astrophysical studies and direct detection experiments. Particularly noteworthy is the comparison with the direct detection limits established by CDEX Jiang et al. (2018), PICO-60 Amole et al. (2019), and XENON1T Aprile et al. (2019) regarding DM-nucleon SD scattering cross-section. This comparison underscores the advantage of constraints derived from local NSs, which extend to DM masses $\sim~{}10^{5}$ GeV.
VI Discussion and Conclusion

### v3

@@ IV Current and Future generation telescopes
Combining data from gamma-ray and neutrino observations enables us to construct a comprehensive picture of DM distribution and behavior across different cosmic scales. Moreover, cross correlating these observations with other astronomical datasets, such as those from gravitational wave detectors and galaxy surveys, enriches our understanding of DM’s role in shaping the cosmos. Multiwavelength studies thus stand as a powerful strategy to decipher the enigmatic properties of DM and elucidate its profound impact on the Universe.
In Sec. III.3, we discuss the shape of gamma-ray and neutrino spectra resulting from DM annihilation through long-lived mediators. In this section, we probe the expected gamma-ray and neutrino flux with the current and future generations of neutrino and gamma-ray telescopes.
@@ IV.1 With current and future neutrino telescopes
Neutrino telescopes offer a novel avenue for probing DM Abbasi et al. (2022a); Miranda et al. (2023) by detecting neutrinos. These telescopes complement traditional detection methods, providing insight into the distribution and properties of DM through neutrino emissions from potential DM sources throughout the Universe.
IceCube, a massive cubic-kilometer neutrino detector situated at the South Pole, was completed in 2010. Originally proposed based on theoretical calculations in 1998, it was envisioned to be sensitive enough to detect high-energy neutrinos from phenomena such as AGN jets or GRBs Waxman and Bahcall (1999). As the first experiment of its kind, IceCube has achieved significant milestones, including the discovery of a diffuse extraterrestrial neutrino flux in 2013 Aartsen et al. (2013) and the detection of neutrino emission from a flaring blazar in 2017 Aartsen et al. (2018a, b). Despite dedicated efforts to identify the sources of cosmic neutrinos through various analyses, such as all-sky survey Aartsen et al. (2020a), transient Abbasi et al. (2022b); Aartsen et al. (2018c, 2020b), and AGN catalog stacking searches Abbasi et al. (2022c); Aartsen et al. (2017a), conclusive results remain elusive. This suggests the possibility of multiple weaker sources contributing to the diffuse flux Murase et al. (2016), such as starburst galaxies Ambrosone et al. (2021), which would necessitate improved pointing resolution Fang et al. (2016).
IceCube Gen 2, an extension of the IceCube Neutrino Observatory in Antarctica, aims to explore the high-energy neutrino sky from TeV to PeV energies Aartsen et al. (2019); Abbasi et al. (2023). It promises five times better sensitivity for neutrino sources compared to the current IceCube detector. With 120 new strings and larger spacing between them, IceCube Gen 2 will enhance sensitivity for neutrinos above 10 TeV and target detection of neutrinos exceeding 100 PeV Abbasi et al. (2023). The ongoing IceCube upgrade will improve detection thresholds down to 1 GeV, enhancing capabilities for various physics studies Abbasi et al. (2021). IceCube Gen 2’s larger instrumented area will enable the detection of fainter neutrino sources, positioning it as the world’s leading neutrino observatory. It holds promise for diverse scientific inquiries, including DM searches, particle physics studies, and the detection of supernova neutrinos Ishihara (2021).
@@ IV.2 With current and future gamma-ray telescopes
Multimessenger astrophysics has become a groundbreaking reality, propelled by recent advancements in gravitational wave, gamma-ray astronomy, and high-energy neutrino detection. These breakthroughs have ushered in a new era of understanding the high-energy astrophysical universe and the mechanisms underlying its most energetic phenomena.
@@
As the next-generation gamma-ray telescope to observe very- to ultrahigh-energy gamma rays, the Southern Wide-field Gamma-ray Observatory (SWGO), a water Cherenkov telescope array, stands at the forefront of this revolutionConceição (2023). Spanning energies from about 30 GeV to a few PeV, SWGO provides a wide field and high-duty cycle view of the southern sky Conceição (2023), complementing existing particle arrays in the Northern Hemisphere like Large High Altitude Air Shower Observatory (LHAASO) LHA (2021). Speaking of which, LHAASO, is a multipurpose facility designed to study cosmic rays and gamma rays across a broad energy spectrum, ranging from sub-TeV to beyond 1 PeV Cao (2010).
@@ V.1 Differential flux sensitivity
Figure 1: Differential flux ($E_{\nu,\gamma}^{2}~{}d\phi_{\nu,\gamma}/dE_{\nu,\gamma}$) sensitivity of all our selected neutrino (left) and gamma-ray (right) detectors.
@@
Finally, for TRIDENT which is expected to be the most sensitive neutrino telescope in the upcoming days, we use the 90% C.L. median sensitivity for the pointlike sources from the expected all-sky survey Ye et al. (2022). In this context, we want to mention that for all neutrino telescopes, we only consider their sensitivity from tracklike events for the sources near declination, $\delta=0^{0}$.
For the gamma-ray telescopes, such as CTA (cta ; Donini et al. (2023)), SWGO (Conceição (2023)), LHAASO (LHA (2021); Conceição (2023)) and Fermi-LAT (fer (2021); Ajello et al. (2021)), we use the differential flux limits for pointlike sources from their respective performance guidelines.
@@ V.2 Variation of differential flux sensitivity of neutrino telescopes with declination
Figure 2: Variation of the differential flux of current and future neutrino telescopes with declination for two DM mass values.
In this section, we study the neutrino signals from NSs resulting from the decay of long-lived mediators produced from DM annihilation in the light of current and future-generation neutrino telescopes. This study specifically focuses on the DM-nucleon scattering cross section in equilibrium, aiming to explore whether NSs offer a promising avenue for probing small DM-nucleon cross sections. Throughout our investigation, we assume that DM number abundance has reached equilibrium, given the age of NSs considered to be higher ($\geq 1$ Myr) than $t_{\rm eq}$. All of our selected neutrino telescopes are equally sensitive for both spin-independent (SI) and spin-dependent (SD) DM-nucleon scattering cross sections as NSs primarily consist of neutrons.
@@
Fig. 2 indicates that for IceCube Aartsen et al. (2020a) and IceCube Gen 2Omeliukh et al. (2021), the highest sensitivity is concentrated around the equator, corresponding to optimal discovery potential, while it is weakest for declination values near sin($\delta$) = ±$\pm 1$ Aartsen et al. (2020a). KM3NeT (with 6 years of projected data) Aiello et al. (2019) and TRIDENT Ye et al. (2022), provide equal sensitivity for all the regions of the sky. This underscores the challenge of IceCube and IceCube Gen 2 of detecting events of interest amidst the background, particularly in regions where a stronger signal is required. For a source located in the southern sky, TRIDENT will have nearly 4 orders of magnitude improvement in sensitivity Ye et al. (2022) compared to IceCube.
Consequently, given the current sensitivity of all these neutrino detectors, TRIDENT has the potential to detect local NSs for DM mass around 1 GeV with $\sigma_{\chi n}$ = $6~{}\times~{}10^{-45}$ cm2. However, for DM mass for $m_{\chi}~{}\sim$ 5 GeV, as shown in Fig. 2, the current and future telescopes are not sufficient unless brighter sources are present or DM-nucleon scattering is enhanced further. The choice of $\sigma_{\chi n}$ = $6~{}\times~{}10^{-45}$ cm2 is motivated by its proximity to the threshold cross section i.e., $\sigma_{\rm{sat}}\simeq 6.58\times 10^{-45}$ cm2 Bramante et al. (2017) of our designated local NSs  [^5] As we use the same mass and radius for each NS, we obtain the same $\sigma_{\rm{sat}}$ value. If $\sigma_{\chi n}$ is less than or equal to $\sigma_{\rm{sat}}$, the DM is expected to lose its kinetic energy and be captured after a single collision. However, once this threshold is exceeded, the effect of multiple scatterings must be considered. To simplify the discussion, in Fig. 2, we focus on the single-scattering scenario..
@@ V.3 Expected bounds on $\sigma_{\chi n}$ from local NSs
V.3.1 Limits from eight local NSs with neutrino telescopes
@@
We consider the DM annihilation process via long-lived mediators as discussed in Sec.III, and DM annihilation into only one specific channel with 100% branching fraction of a species (neutrino or gamma-ray). As a result, we obtain the maximum upper limit on the capture rate which is then translated into the DM scattering cross section.
Figure 3, with conservative measures, we translate the differential flux sensitivity (Fig. 2) at $\delta=0^{0}$, of our selected neutrino detectors at Sec. V.2 and derive the stacked constraints [^6] For obtaining the stacked flux values, we combine the expected flux from each NS. This treatment generally applies to improve the sensitivity, especially for the null or faint detection. on $\sigma_{\chi n}$ as a function of the DM mass, $m_{\chi}$ by combining the flux limits from 8 local NSs (Eqs. (10) and (11)). The limits shown in Fig. 3 emerge from the intricate interplay of the DM capture rate, DM density profile, and DM mass range and serve as the current constraints on $\sigma_{\chi n}$ for 8 nearby local NSs. We derive the stacked limits for the NFW density profile as mentioned in Sec. III.1. To ensure the robustness of our findings, we validate the results against those obtained for the cored Burkert profile Burkert (1995); Salucci and Burkert (2000). However, our analysis reveals $\sigma_{\chi n}$ has no significant dependence on the density profile for local NSs.
Figure3 highlights that with 8 local NSs, we only obtain the limits for DM mass $<$ 1 GeV, and only IceCube Gen 2 and TRIDENT can provide any meaningful bounds. Here we want to stress that as for Fig.3, we currently get the limits only for DM mass $<$ 1 GeV, we correct the DM capture rate by implementing the Pauli Blocking effect (Eq. (7)). Furthermore, in Fig. 3, we compare the limits with experimental bounds obtained from direct detection bounds, such as CRESST-IIIAngloher et al. (2024) with a total exposure of 3.64 kg-day, CDEX10Liu et al. (2022) with 205.4 kg-day of exposure and XENON1TAprile et al. (2019a) for one tonne-year of exposure. For DM mass $<$ 1 GeV, a handful of nearby NSs provide stronger bounds than obtained from XENON1T Aprile et al. (2019a).
The observations of Fig.3 is consistent with the finding of Fig. 2. An NS with a certain maximum capture rate will have an emission of maximum flux (neutrino or gamma-ray) due to the annihilation of DM. If the DM annihilation flux from the local NS remains below the sensitivity of the detector, it will remain undetected and no constraints on DM mass and its scattering cross section can be placed.
V.3.2 Limits from the local population of NSs with neutrino and gamma-ray telescopes
@@
Figure 4: DM annihilation spectra through long-lived mediators for three scattering cross section values expected from the local NS population with $N_{\rm NS}~{}\approx~{}10^{5}$ within 0.6 kpc from Earth.
In Sec. V.3.1, we observe that the set of local NSs considered in Sec.II.1 fails to provide any conclusive limit upon DM capture for GeV scale DM. This also indicates that the bound obtained from only a few locally detected NSs is less significant from the statistical point of view and also less detectable with the present sensitivity of the detectors. However, the population study of local NSs with a distribution based on the DM halo profile and CGW searches has the potential to improve the present significantly reported limits on DM capture rate for GeV scale DM.
Following the population model proposed by Ref. Reed et al. (2021), we anticipate $N_{\rm NS}~{}\approx~{}10^{5}$ within 0.6 kpc from Earth for NS ellipticity $\epsilon~{}\sim~{}10^{-5}$. In Fig. 4, we present the expected differential gamma-ray and neutrino flux (horizontal lines) from $10^{5}$ the number of local NSs for the NFW halo profile. We use Eqs. (10)-(11) to obtain the expected flux for energy ranging between $1-10^{6}$ GeV for three scattering cross section ($\sigma_{\chi n}$) values. For a chosen value of $\sigma_{\chi n}$, both gamma-ray and neutrino spectra result in the same flux limits due to box-shaped annihilation spectra. We further compare the upper limit of differential flux with the current and projected sensitivities of the gamma-ray and neutrino detectors (from Fig. 1) as shown in Fig. 4. We observe that CTA, LHASSO and SGWO is capable of probing DM annihilation flux from the local NS population for $\sigma_{\chi n}\leq 10^{-45}$ cm2 in the mass range (or energy range, for boxlike spectra, they are equivalent) $m_{\chi}\geq 100$ GeV while Fermi-LAT and TRIDENT can probe for $m_{\chi}\leq 100$ GeV. The projected sensitivity of IceCube Gen 2 is promising for probing $m_{\chi}\geq 10^{5}$ GeV for $\sigma_{\chi n}~{}=~{}10^{-44}-10^{-45}$ cm2. Moreover, we also find that IceCube and KM3Net are not sufficiently sensitive for $\sigma_{\chi n}~{}\leq~{}10^{-44}$ cm2. In conclusion, Fig. 4 further indicates that although most experimental limits are sensitive to detection of a high annihilation flux with $\sigma_{\chi n}\geq 10^{-45}$ cm2, a lower DM scattering cross section $\sigma_{\chi n}<10^{-46}$ cm2 is unreachable with the current detector sensitivity with respect to the upper limit for $10^{5}$ number of local NS population from CGW searches.
@@
Figure 5: Comprehensive limits on $\sigma_{\chi n}$ from the local NS population with CGW searches (solid lines) and comparison with several DM direct detection experiments and limits obtained from astrophysical studies.
In Fig. 5, we convert the differential flux sensitivity from Fig. 1 to show the projected bounds for $10^{5}$ local NSs. In Fig. 3, due to the limited statistics from just eight NSs, we adopt a conservative approach by using the differential flux sensitivity from Fig. 2 at $\delta=0^{\circ}$. For Fig. 5, where we expect to improve our bounds with the larger local NS population, we use the detector sensitivity from Fig. 1, which varies as a function of energy. Comparing with Fig. 3, we find that the bounds predicted by the CGW search with the local NS population provide more stringent limits on the DM scattering cross section across a wide range of DM mass, from a few GeV to $10^{6}$ GeV. We restrict ourselves to the DM mass, $10^{6}$ GeV to minimize the multiple scattering aspects of DM with nucleon for the heavy DM mass Bell et al. (2020).
In Fig. 5, we further extend our analysis to compare our findings with the results obtained from various direct detection experiments. Specifically, we consider CDEX10Liu et al. (2022) with 205.4 kg-day, CRESST-IIIAngloher et al. (2024) with a total exposure of 3.64 kg-day, LZAalbers et al. (2023) with 60 live days, PandaX-IIXia et al. (2019) with a total exposure of 54-ton-days, and XENON1TAprile et al. (2019a, b) using one tonne-year of exposure corresponds to DM-neutron SD scattering cross section. We also compare our results with astrophysical limits from IceCube for Sun Abbasi et al. (2022a); Aguilar Sánchez, Juan Antonio (2019) and indirect search bounds from H.E.S.S. for Galactic population of NS Leane et al. (2021). This comparison underscores the advantage of constraints derived from local NSs, which extend to DM masses $\sim~{}10^{6}$ GeV.
From Fig. 5, we observe that at low mass regime (between 1-100 GeV), only Fermi-LAT and TRIDENT provide the bounds on $\sigma_{\chi n}$ based on their projected sensitivity, with TRIDENT imposing the strongest limit, i.e., $\sigma_{\chi n}~{}\sim~{}10^{-45}$ cm2, which is several orders of magnitude stronger than the present direct detection and astrophysical bounds. Similarly, in the heavy DM regime ($m_{\chi}\geq 1000$ GeV), we obtain similar bounds on $\sigma_{\chi n}$ from searches of the local population of NSs with the projected sensitivity of CTA, IceCube Gen 2, and LHAASO, while SWGO imposes the most stringent limit of $\sigma_{\chi n}~{}\sim~{}10^{-45}$ cm2, surpassing direct detection experiments by a considerable margin. Among all astrophysical bounds compared in Fig. 5, H.E.S.S. bound from Galactic NS contribution is on par with our limits on $\sigma_{\chi n}$. Therefore, Fig. 5 suggests that CGW missions with improved search capabilities for local NSs have the potential to tighten the limits on the DM scattering cross section across a wide range of DM masses, which we are generally unable to probe with a few detected local NSs.
VI Conclusion